Herschall has $18,000 to invest. He invest 2/3 in one account and the rest in a second account. The first returned 2% more than the second account. The total return on his investment was $1050. What were the percentage rates of the two accounts.
Answer by stanbon(57323)

Herschall has $18,000 to invest. He invest 2/3 in one account and the rest in a second account.
----
(2/3)18000 = 12,000
(1/3)18000 = 6,000
----
The first returned 2% more than the second account. The total return on his investment was $1050. What were the percentage rates of the two accounts.
-------------
Equation:
interest + interest = 1050
12000x + 6000(x+0.02) = 1050
---
12000x + 6000x + 120 = 1050
18000x + 120 = 1050
18000x = 930
x = 0.0517 = 5.17% (rate on the 12000)
x+0.02 = 7.17% (rate on the 6000)
=====================================
Cheers,
Stan H.
================



========================